# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ 6 Sufficient conditions for (11)
Proof of Proposition 6.6.
Recall that, for $1\leq r\leq d$, the vectors $a_{1},\ldots,a_{r}$ of $\mathbb{R}^{d}$ are linearly dependent if and only if every $r\times r$ submatrix of $\mathbb{A}:=(a_{1},\ldots,a_{r})$ has determinant equal to zero. Let $\mathcal{J}_{r}$ be the collection of subsets of $\{1,\ldots,d\}$ with cardinality equal to $r$. For $J\in\mathcal{J}_{r}$, $J=\{j_{1},\ldots j_{r}\}$, let $p_{J}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$ be the projection $(x_{1},\ldots,x_{d})^{\top}\mapsto(x_{j_{1}},\ldots,x_{j_{r}})^{\top}$. Clearly, for every $J\in\mathcal{J}_{r}$, $\widetilde{p}_{J}(\mathbb{A}):=(p_{J}(a_{1}),\ldots,p_{J}(a_{r}))$ is an $r\times r$ submatrix of $\mathbb{A}$, and every $r\times r$ submatrix of $\mathbb{A}$ arises in this way for some $J\in\mathcal{J}_{r}$. Note that $a_{1},\ldots,a_{r}$ are linearly dependent if and only if $f(\mathbb{A})=0$ where $f:\mathbb{R}^{d\times r}\rightarrow[0,\infty)$ is defined by
@@
It is easily realized that $f$ is a continuous function (as composition of continuous functions), and so $f$ is measurable. Set $\mathbb{X}=(X_{1},\ldots,X_{r})$ and
$E:=\{\text{the random vectors $X_{1},\ldots,X_{r}$ are linearly independent}\}.$
We have
$\mathbb{P}(E)=1-\mathbb{P}(f(\mathbb{X})=0)\geq 1-\mathbb{P}({\rm det}(\widetilde{p}_{\{1,\ldots,r\}}(\mathbb{X}))=0).$
Since ${\rm det}(\widetilde{p}_{\{1,\ldots,r\}}(\cdot))$ is a polynomial function, by the absolute continuity of the law of ${\rm vec}((X_{1},\ldots,X_{r}))$ with respect to the Lebesgue measure we have
@@ 7 Numerical illustrations
In this section we provide numerical illustrations of Theorem 5.1. More precisely, we consider two specific models of neural networks with dependent and heavy-tailed weights, we sample from the posterior distribution of the output at a finite and at an infinite width and we verify the convergence, as the number of nodes in the hidden layers grows large. We start introducing the models.
7.1 Model 1
The first model that we analyze is a neural network with dependent weights, as defined in Sections 2.1 and 2.2, with $C_{B}>0$ and $V_{n_{\ell},j}^{(\ell)}:=Y_{j}^{(\ell)}/n_{\ell}$, $j=1,\ldots,n_{\ell}$, $\ell=1,\ldots,L$, where, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell=1,\ldots,L}$ are independent random variables and $\{Y_{j}^{(\ell)}\}_{j\geq 1}$ are identically distributed with $\mathbb{P}_{{\rm prior}}(Y_{1}^{(\ell)}>0)=1$ and $\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)}\in(0,\infty)$. By the law of the large numbers, for each $\ell=1,\ldots,L$, we have
$\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)},\quad\text{$\mathbb{P}_{{\rm prior}}$-a.s., as $n_{\ell}\to\infty$.}$
Therefore, the assumption $(ii)$ of Theorem 4.1 holds with $a^{(\ell)}=\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)}$ and $\rho^{(\ell)}\equiv 0$. If the inputs are e.g. distinct and the activation function $\sigma$ is e.g. Lipschitz continuous and nonlinear, then all the hypotheses of Theorem 6.2 are fullfilled and the corresponding assumption (11) holds.
To make the model more realistic, under the prior, the weights should be heavy-tailed (see the introduction and the references [29, 38]). For instance, this happens if, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell=1,\ldots,L}$ are independent and identically distributed random variables with $Y_{1}^{(1)}\stackrel{{\scriptstyle\mathcal{L}}}{{=}}(WE)^{2}$, where $WE$ has the Weibull distribution with parameters $(1,1/2)$, i.e., it has density (with respect to the Lebesgue measure)
$f_{WE}(x):=\frac{1}{2}x^{-1/2}\mathrm{e}^{-\sqrt{x}},\quad x>0$
@@ 7.1 Model 1
(we could work with other heavy-tailed distributions, however, since we are interested in simulating the model, we prefer to make a specific choice of the parameters soon at this stage). Indeed, in such a case, the Laplace transform of $W_{11}^{(2)}$ on $\mathbb{R}\setminus\{0\}$ is equal to infinity, as a simple calculation shows. Note also that a standard computation gives $\mathbb{E}_{{\rm prior}}(WE)^{2}=24$, and so in this specific case $a^{(\ell)}=24$, $\ell=1,\ldots,L$.
7.2 Model 2
@@
The second model that we analyze is a neural network with dependent weights, as defined in Sections 2.1 and 2.2, with $C_{B}>0$ and $V_{n_{\ell},j}^{(\ell)}:=\frac{\pi^{2}}{n_{\ell}^{2}}Y_{j}^{(\ell)}$, $j=1,\ldots,n_{\ell}$, $\ell=1,\ldots,L$, where, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell=1,\ldots,L}$ is a family of independent and identically distributed random variables with $Y_{1}^{(1)}\stackrel{{\scriptstyle d}}{{=}}(HC)^{2}$, where $HC$ denotes a random variable distributed according to the half-Cauchy law, i.e., with probability density (with respect to the Lebesgue measure)
$f_{HC}(x):=\frac{2}{\pi(1+x^{2})}\mathbb{1}\{x>0\}.$
Note that $\mathbb{P}_{{\rm prior}}(Y_{1}^{(1)}>0)=1$, $\ell=1,\ldots,L$. Under $\mathbb{P}_{{\rm prior}}$, it turns out (see Appendix E in [26]) that, for $\ell=1,\ldots,L$,
$\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to{\rm ID}(0,\rho),\quad\text{in distribution, as $n_{\ell}\to\infty$,}$
where
$\rho(\mathrm{d}x):=x^{-3/2}\mathbb{1}\{x>0\}\mathrm{d}x.$ (36)
Therefore, the assumption $(ii)$ of Theorem 4.1 holds with $a^{(\ell)}=0$ and $\rho^{(\ell)}=\rho$ for each $\ell$. Note that, for any $\varepsilon>0$,
$\rho((\varepsilon,\infty))=2\varepsilon^{-1/2}\quad\text{and}\quad\rho((0,\varepsilon])=+\infty,$
@@
and therefore a Poisson process with mean measure $\rho$ have infinitely many points on $(0,\infty)$. Consequently, the assumption (17) holds. If the inputs are such that (18) is satisfied and the activation function $\sigma$ is Lipschitz and non constant, then all the hypotheses of Theorem 6.1 are fullfilled and the corresponding assumption (11) holds.
Here again, to make the model more realistic, under the prior, the weights should be heavy-tailed (see the introduction and the references [29, 38]). This is the case for the Model 2. Indeed, the Laplace transform of $W_{11}^{(2)}$ on $\mathbb{R}\setminus\{0\}$ is equal to infinity, as can be easily verified by a simple computation.
7.3 Simulation of Model 1
@@
We consider the Model 1 with $C_{B}:=1$, $C_{W}:=1$, $L:=2$, $n_{0}:=4$, $n_{3}:=1$, $n_{1}=n_{2}=n\in\{2,4,8,16,32\}$, $x(1):=(1,0,0,0)^{\top}$, $x(2):=(0,1,0,0)^{\top}$, $x(3):=(0,0,1,0)^{\top}$, $d:=3$, $y(1):=f(x(1))$, $y(2):=f(x(2))$, $y(3):=f(x(3))$, where $f(v_{1},v_{2},v_{3},v_{4}):=v_{1}^{2}+2v_{1}v_{2}+v_{3}^{2}$, and activation function $\sigma(x):=\max\{0,x\}$. Then
$\mathbb{K}^{(1)}(\mathbb{x}):=(K^{(1)}(x(i),x(i^{\prime})))_{1\leq i,i^{\prime}\leq 3},\qquad\text{where}\qquad K^{(1)}(x(i),x(i^{\prime})):=1+\frac{1}{4}x(i)^{\top}x(i^{\prime}),$
@@
By Theorem 5.1 if we denote by $G^{(3)}(\mathbb{x})$ the infinite-width limit of the posterior law of the output, then
$G^{(3)}(\mathbb{x})\sim\mathcal{N}_{3}(\lambda^{(3)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(3)}(\mathbb{x})).$
Since, for $A\in\mathcal{B}(\mathbb{R}^{3})$,
$\mathbb{P}_{{\rm posterior}}(Z^{(3)}(\mathbb{x})\in A)=\mathfrak{c}\mathbb{E}_{{\rm prior}}\left[\mathbb{1}_{A}(Z^{(3)}(\mathbb{x}))\exp\left(-\sum_{i=1}^{3}\|Z^{(3)}(x(i))-y(i)\|^{2}\right)\right],$
@@
if the number of nodes is kept fixed, we sample from the posterior law of $Z^{(3)}(\mathbb{x})$ simulating this random variable under the prior, and then using a Monte Carlo estimator. See Figure 1 for a numerical validation of Theorem 5.1.
@@
Figure 1: Simulation of Model 1. The non-green curves are estimates of the one dimensional marginal distribution functions (with $n_{1}=n$ nodes in first hidden layer and $n_{2}=2n$ in second hidden layer, for different values of $n=4,8,16,32$) of the distribution function of the posterior law of the 33-variate output $Z^{(3)}(\mathbb{x})$. The green curves are the one dimensional marginal distribution functions of the distribution function of the 33-variate infinite-width limit $G^{(3)}(\mathbb{x})$. All the parameters are specified in Section 7.3.
7.4 Simulation of Model 2
We consider the Model 2 with $C_{B}:=1$, $C_{W}:=1$, $L=1$, $n_{0}:=4$, $n_{1}\in\{2,4,8,16,32\}$, $n_{2}:=1$, $d:=3$, $x(1):=(1,0,0,0)^{\top}$, $x(2):=(0,1,0,0)^{\top}$, $x(3):=(0,0,1,0)^{\top}$, $y(1):=f(x(1))$, $y(2):=f(x(2))$, $y(3):=f(x(3))$, where $f(v_{1},v_{2},v_{3},v_{4})=v_{1}^{2}+2v_{1}v_{2}+v_{3}^{2}$, and activation function $\sigma(x):=\max\{0,x\}$. Then
$\mathbb{K}^{(1)}(\mathbb{x}):=(K^{(1)}(x(i),x(i^{\prime})))_{1\leq i,i^{\prime}\leq 3},\qquad\text{where}\qquad K^{(1)}(x(i),x(i^{\prime})):=1+\frac{1}{4}x(i)^{\top}x(i^{\prime}),$
and
$\mathbb{K}^{(2)}(\mathbb{x}):=1_{3}1_{3}^{\top}+\sum_{j=1}^{\infty}T_{j}^{(1)}\sigma(\zeta_{j}^{(1)}(\mathbb{x}))\sigma(\zeta_{j}^{(1)}(\mathbb{x}))^{\top}.$
@@ 7.4 Simulation of Model 2
Here, $\{\zeta_{j}^{(1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of i.i.d. random vectors with $\zeta_{1}^{(1)}(\mathbb{x})\sim\mathcal{N}_{3}(0,\mathbb{K}^{(1)}(\mathbb{x}))$, independent of $\{T_{j}^{(1)}\}_{j\geq 1}$, which are the points of a Poisson process on $(0,\infty)$ with mean measure $\rho$ defined by (36).
By Theorem 5.1 if we denote by $G^{(2)}(\mathbb{x})$ the infinite-width limit of the posterior law of the output, it holds
$G^{(2)}(\mathbb{x})\sim\mathcal{GM}_{3}^{\mathbb{S}}(\lambda^{(2)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(2)}(\mathbb{x})).$
@@
Figure 2: Simulation of Model 2. The non-green curves are estimates of the one dimensional marginal distribution functions (for different values $n_{1}=n=2,4,8,16,32$ of the number of nodes in the hidden layer) of the posterior law of the 33-variate output $Z^{(2)}(\mathbb{x})$. The green curves are the one dimensional marginal distribution functions of the distribution function of the 33-variate infinite-width limit $G^{(2)}(\mathbb{x})$. All the parameters are specified in Section 7.4.
@@
$\quad=\mathfrak{c}\mathbb{E}_{{\rm prior}}\left[\frac{\exp\left(\mathrm{Tr}(\mathbb{y}(\mathbb{Id}_{3}+(2\mathbb{K}^{(2)}(\mathbb{x}))^{-1})^{-1})\mathbb{y}^{\top}\right)}{(\mathrm{det}(\mathbb{Id}_{3}+2\mathbb{K}^{(2)}(\mathbb{x})))^{n_{2}/2}}\int_{A}\varphi_{(\lambda^{(2)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(2)}(\mathbb{x}))}(\xi)\mathrm{d}\xi\right].$
@@
and, as noticed in Appendix E.3.2 of [26],
$T_{(j)}^{(1)}=\frac{4}{(\sum_{k=1}^{j}E_{k})^{2}},$
where $\{E_{k}\}_{k\geq 1}$ is a sequence of independent random variables with the exponential law with mean $1$.
If the number of nodes is kept fixed, we sample from the posterior law of $Z^{(\ell)}(\mathbb{x})$, $\ell=1,2$, simulating these random variables under the prior, and then using a Monte Carlo estimator as described in Section 7.3.
See Figure 2 for another numerical validation of Theorem 5.1.
Appendix: Proof of Theorem 4.2
The proof of Theorem 4.2 is based on two lemmas. The first one, Lemma 7.1, provides an alternative representation of the neural network; its proof is omitted since it is similar to the proof of Lemma 3 in [27]. The second one, Lemma 7.2, is concerned with the weak limit of triangular arrays; its proof can be found in [26], see Corollary 41 and the last lines of the proof of Theorem 16 therein.
Let $\{\widetilde{N}_{jr}^{(\ell)}\}_{j,\ell\geq 1;r=1,\ldots,d}$ be a family of independent standard normal random variables, independent of the family $\{V_{n_{\ell-1},j}^{(\ell-1)}\}_{\ell=1,\ldots,L+1;j=1,\ldots n_{\ell-1}}$. For $\ell=1,\ldots,L$, we define the symmetric $d\times d$ positive semi-definite random matrices
$\mathbb{H}^{(\ell)}(\mathbb{x})=C_{B}1_{d}1_{d}^{\top}+C_{W}\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\sigma(\gamma_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\gamma_{j}^{(\ell-1)}(\mathbb{x}))^{\top}$ (37)
and
$\mathbb{H}^{(0)}(\mathbb{x})=\mathbb{K}^{(1)}(\mathbb{x}):=C_{B}+\frac{C_{W}}{n_{0}}\mathbb{x}^{\top}\mathbb{x}.$
Here, for $\ell=1,\ldots,L$ and $j=1,\ldots,n_{\ell}$,

### v5

@@ 6.2. Proof of Propositions 1, 2, 3 and 4
Proof of Proposition 4.
Recall that, for $r\in[q]$, the vectors $a_{1},\ldots,a_{r}$ of $\mathbb{R}^{q}$ are linearly dependent if and only if every $r\times r$ submatrix of $\mathbb{A}:=(a_{1},\ldots,a_{r})$ has determinant equal to zero. Let $\mathcal{J}_{r}$ be the collection of subsets of $[q]$ with cardinality equal to $r$. For $J\in\mathcal{J}_{r}$, $J=\{j_{1},\ldots j_{r}\}$, let $p_{J}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{r}$ be the projection $(x_{1},\ldots,x_{q})^{\top}\mapsto(x_{j_{1}},\ldots,x_{j_{r}})^{\top}$. Clearly, for every $J\in\mathcal{J}_{r}$, $\widetilde{p}_{J}(\mathbb{A}):=(p_{J}(a_{1}),\ldots,p_{J}(a_{r}))$ is an $r\times r$ submatrix of $\mathbb{A}$, and every $r\times r$ submatrix of $\mathbb{A}$ arises in this way for some $J\in\mathcal{J}_{r}$. Note that $a_{1},\ldots,a_{r}$ are linearly dependent if and only if $f(\mathbb{A})=0$ where $f:\mathbb{R}^{q\times r}\rightarrow[0,\infty)$ is defined by
@@
It is easily realized that $f$ is a continuous function (as composition of continuous functions), and so $f$ is measurable. Set $\mathbb{X}=(X_{1},\ldots,X_{r})$ and $E:=\{\text{the r.v.'s $X_{1},\ldots,X_{r}$ are linearly independent}\}.$ We have $\mathbb{P}(E)=1-\mathbb{P}(f(\mathbb{X})=0)\geq 1-\mathbb{P}({\rm det}(\widetilde{p}_{\{1,\ldots,r\}}(\mathbb{X}))=0).$ Since ${\rm det}(\widetilde{p}_{\{1,\ldots,r\}}(\cdot))$ is a polynomial function, by the absolute continuity of the law of ${\rm vec}((X_{1},\ldots,X_{r}))$ with respect to the Lebesgue measure we have
@@ 7. Numerical illustrations
In this section we provide numerical illustrations of Theorem 3. More precisely, we consider two models of neural networks with dependent and heavy-tailed weights, we sample from the posterior distribution of the output at a finite and at an infinite width and we verify the convergence, as the number of nodes in the hidden layers grows large. We start introducing the models.
7.1. Model 1
@@
The first model that we analyze is a neural network with dependent weights, as defined in Sections 2.1 and 2.2, with $C_{B}>0$ and $V_{n_{\ell},j}^{(\ell)}:=Y_{j}^{(\ell)}/n_{\ell}$, $j\in[n_{\ell}]$, $\ell\in[L]$, where, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell\in[L]}$ are independent r.v.’s and $\{Y_{j}^{(\ell)}\}_{j\geq 1}$ are identically distributed with $\mathbb{P}_{{\rm prior}}(Y_{1}^{(\ell)}>0)=1$ and $\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)}\in(0,\infty)$. By the law of the large numbers, for each $\ell\in[L]$, we have $\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)},$ $\mathbb{P}_{{\rm prior}}$-a.s., as $n_{\ell}\to\infty$. Therefore, the assumption $(ii)$ of Theorem 1 holds with $a^{(\ell)}=\mathbb{E}_{{\rm prior}}Y_{1}^{(\ell)}$ and $\rho^{(\ell)}\equiv 0$. If the inputs are e.g. distinct and the activation function $\sigma$ is e.g. Lipschitz continuous and nonlinear, then all the hypotheses of Theorem 5 are fullfilled and the corresponding assumption (5.2) holds. To make the model more realistic, under the prior, the weights should be heavy-tailed (see the introduction and the references [29, 37]). For instance, this happens if, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell\in[L]}$ are i.i.d. r.v.’s with $Y_{1}^{(1)}\stackrel{{\scriptstyle\mathcal{L}}}{{=}}(WE)^{2}$, where $WE$ has the Weibull distribution with parameters $(1,1/2)$, i.e., it has density (with respect to the Lebesgue measure) $f_{WE}(x):=\frac{1}{2}x^{-1/2}\mathrm{e}^{-\sqrt{x}},\quad x>0$ (we could work with other heavy-tailed distributions, however, since we are interested in simulating the model, we prefer to make a specific choice of the parameters soon at this stage). Indeed, in such a case, the Laplace transform of $W_{11}^{(2)}$ on $\mathbb{R}\setminus\{0\}$ is equal to infinity, as a simple calculation shows. Note also that a standard computation gives $\mathbb{E}_{{\rm prior}}(WE)^{2}=24$, and so in this specific case $a^{(\ell)}=24$, $\ell\in[L]$.
7.2. Model 2
@@
The second model that we analyze is a neural network with dependent weights, as defined in Sections 2.1 and 2.2, with $C_{B}>0$ and $V_{n_{\ell},j}^{(\ell)}:=\frac{\pi^{2}}{n_{\ell}^{2}}Y_{j}^{(\ell)}$, $j\in[n_{\ell}]$, $\ell\in[L]$, where, under $\mathbb{P}_{{\rm prior}}$, $\{Y_{j}^{(\ell)}\}_{j\geq 1,\ell\in[L]}$ is a family of i.i.d. r.v.’s with $Y_{1}^{(1)}\stackrel{{\scriptstyle d}}{{=}}(HC)^{2}$, where $HC$ denotes a r.v. distributed according to the half-Cauchy law, i.e., with probability density (with respect to the Lebesgue measure) $f_{HC}(x):=\frac{2}{\pi(1+x^{2})}\mathbb{1}\{x>0\}.$ Note that $\mathbb{P}_{{\rm prior}}(Y_{1}^{(1)}>0)=1$, $\ell\in[L]$. Under $\mathbb{P}_{{\rm prior}}$, it turns out (see Appendix E in [26]) that, for $\ell\in[L]$, $\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to{\rm ID}(0,\rho),$ in distribution, as $n_{\ell}\to\infty$, where
(7.1) $\rho(\mathrm{d}x):=x^{-3/2}\mathbb{1}\{x>0\}\mathrm{d}x.$
@@
Therefore, the assumption $(ii)$ of Theorem 1 holds with $a^{(\ell)}=0$ and $\rho^{(\ell)}=\rho$ for each $\ell$. Note that, for any $\varepsilon>0$, $\rho((\varepsilon,\infty))=2\varepsilon^{-1/2}$ and $\rho((0,\varepsilon])=+\infty,$ and therefore a Poisson process with mean measure $\rho$ have infinitely many points on $(0,\infty)$. Consequently, the assumption (6.1) holds. If the inputs are such that (6.2) is satisfied and the activation function $\sigma$ is Lipschitz and non constant, then all the hypotheses of Theorem 4 are fullfilled and the corresponding assumption (5.2) holds. Here again, to make the model more realistic, under the prior, the weights should be heavy-tailed, see the introduction and [29, 37]. This is the case for the Model 2. Indeed, the Laplace transform of $W_{11}^{(2)}$ on $\mathbb{R}\setminus\{0\}$ is equal to infinity, as can be easily verified by a simple computation.
7.3. Simulation of Model 1
@@
We consider the Model 1 with $C_{B}:=1$, $C_{W}:=1$, $L:=2$, $n_{0}:=4$, $n_{3}:=1$, $n_{1}=n\in\{2,4,8,16,32\}$, $n_{2}=2n$, $d:=3$, $x(1):=(1,0,0,0)^{\top}$, $x(2):=(0,1,0,0)^{\top}$, $x(3):=(0,0,1,0)^{\top}$, $d:=3$, $y(1):=f(x(1))$, $y(2):=f(x(2))$, $y(3):=f(x(3))$, where $f(v_{1},v_{2},v_{3},v_{4}):=v_{1}^{2}+2v_{1}v_{2}+v_{3}^{2}$, and activation function $\sigma(x):=\max\{0,x\}$. Then
(7.2) $\mathbb{K}^{(1)}(\mathbb{x}):=(K^{(1)}(x(i),x(i^{\prime})))_{1\leq i,i^{\prime}\leq 3},\qquad\text{where}\qquad K^{(1)}(x(i),x(i^{\prime})):=1+\frac{1}{4}x(i)^{\top}x(i^{\prime}),$
@@
By Theorem 3 if we denote by $G^{(3)}(\mathbb{x})$ the infinite-width limit of the posterior law of the output, then $G^{(3)}(\mathbb{x})\sim\mathcal{N}_{3}(\lambda^{(3)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(3)}(\mathbb{x})).$ Since, for $A\in\mathcal{B}(\mathbb{R}^{3})$,
$\mathbb{P}_{{\rm posterior}}(Z^{(3)}(\mathbb{x})\in A)\propto\mathbb{E}_{{\rm prior}}\left[\mathbb{1}_{A}(Z^{(3)}(\mathbb{x}))\exp\left(-\sum_{i=1}^{3}\|Z^{(3)}(x(i))-y(i)\|^{2}\right)\right],$
@@
if the number of nodes is kept fixed, we sample from the posterior law of $Z^{(3)}(\mathbb{x})$ simulating this random variable under the prior, and then using a Monte Carlo estimator. See Figure 1 for a numerical validation of Theorem 3.
@@
Figure 1. Simulation of Model 1. The non-green curves are estimates of the one dimensional marginal distribution functions (with $n_{1}=n$ nodes in the first hidden layer and $n_{2}=2n$ in the second hidden layer, for different values of $n=4,8,16,32$) of the distribution function of the posterior law of the 33-variate output $Z^{(3)}(\mathbb{x})$. The green curves are the one dimensional marginal distribution functions of the distribution function of the 33-variate infinite-width limit $G^{(3)}(\mathbb{x})$. All the other parameters of the model are specified in Section 7.3.
7.4. Simulation of Model 2
@@
We consider the Model 2 with $C_{B}:=1$, $C_{W}:=1$, $L=1$, $n_{0}:=4$, $n_{1}\in\{2,4,8,16,32\}$, $n_{2}:=1$, $d:=3$, $x(1):=(1,0,0,0)^{\top}$, $x(2):=(0,1,0,0)^{\top}$, $x(3):=(0,0,1,0)^{\top}$, $y(1):=f(x(1))$, $y(2):=f(x(2))$, $y(3):=f(x(3))$, where $f(v_{1},v_{2},v_{3},v_{4})=v_{1}^{2}+2v_{1}v_{2}+v_{3}^{2}$, and activation function $\sigma(x):=\max\{0,x\}$. Then $\mathbb{K}^{(1)}(\mathbb{x})$ is given by (7.2) and $\mathbb{K}^{(2)}(\mathbb{x}):=1_{3}1_{3}^{\top}+\sum_{j=1}^{\infty}T_{j}^{(1)}\sigma(\zeta_{j}^{(1)}(\mathbb{x}))\sigma(\zeta_{j}^{(1)}(\mathbb{x}))^{\top}.$ Here, $\{\zeta_{j}^{(1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of i.i.d. r.v.’s with $\zeta_{1}^{(1)}(\mathbb{x})\sim\mathcal{N}_{3}(0,\mathbb{K}^{(1)}(\mathbb{x}))$, independent of $\{T_{j}^{(1)}\}_{j\geq 1}$, which are the points of a Poisson process on $(0,\infty)$ with mean measure $\rho$ defined by (7.1). By Theorem 3 if we denote by $G^{(2)}(\mathbb{x})$ the infinite-width limit of the posterior law of the output, it holds $G^{(2)}(\mathbb{x})\sim\mathcal{GM}_{3}^{\mathbb{S}}(\lambda^{(2)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(2)}(\mathbb{x})).$
@@
Figure 2. Simulation of Model 2. The non-green curves are estimates of the one dimensional marginal distribution functions (for different values $n_{1}=n=2,4,8,16,32$ of the number of nodes in the hidden layer) of the posterior law of the 33-variate output $Z^{(2)}(\mathbb{x})$. The green curves are the one dimensional marginal distribution functions of the distribution function of the 33-variate infinite-width limit $G^{(2)}(\mathbb{x})$. All the other parameters of the model are specified in Section 7.4.
@@
$\quad\propto\mathbb{E}_{{\rm prior}}\left[\frac{\exp\left(\mathrm{Tr}(\mathbb{y}(\mathbb{Id}_{3}+(2\mathbb{K}^{(2)}(\mathbb{x}))^{-1})^{-1})\mathbb{y}^{\top}\right)}{(\mathrm{det}(\mathbb{Id}_{3}+2\mathbb{K}^{(2)}(\mathbb{x})))^{n_{2}/2}}\int_{A}\varphi_{(\lambda^{(2)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(2)}(\mathbb{x}))}(\xi)\mathrm{d}\xi\right].$
@@
and, as noticed in Appendix E.3.2 of [26], $T_{(j)}^{(1)}=\frac{4}{(\sum_{k=1}^{j}E_{k})^{2}},$ where $\{E_{k}\}_{k\geq 1}$ is a sequence of i.i.d. r.v.’s with the exponential law with mean $1$. If the number of nodes is kept fixed, we sample from the posterior law of $Z^{(\ell)}(\mathbb{x})$, $\ell=1,2$, simulating these r.v.’s under the prior, and then using a Monte Carlo estimator as described in Section 7.3. See Figure 2 for another numerical validation of Theorem 3.
Appendix A Proof of Theorem 2
The proof of Theorem 2 is based on two lemmas. The first one, Lemma 3, provides an alternative representation of the neural network; its proof is omitted since it is similar to the proof of Lemma 3 in [27]. The second one, Lemma 4, is concerned with the weak limit of triangular arrays; its proof can be found in [26], see Corollary 41 and the last lines of the proof of Theorem 16 therein.
Let $\{\widetilde{N}_{jr}^{(\ell)}\}_{j,\ell\geq 1;r\in[d]}$ be a family of i.i.d. standard normal r.v.’s, independent of the family $\{V_{n_{\ell-1},j}^{(\ell-1)}\}_{\ell\in[L+1];j\in[n_{\ell-1}]}$. For $\ell\in[L]$, we define the symmetric $d\times d$ positive semi-definite random matrices
(A.1) $\mathbb{H}^{(\ell)}(\mathbb{x})=C_{B}1_{d}1_{d}^{\top}+C_{W}\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\sigma(\gamma_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\gamma_{j}^{(\ell-1)}(\mathbb{x}))^{\top}$
and $\mathbb{H}^{(0)}(\mathbb{x})=\mathbb{K}^{(1)}(\mathbb{x}):=C_{B}+\frac{C_{W}}{n_{0}}\mathbb{x}^{\top}\mathbb{x}.$ Here, for $\ell\in[L]$ and $j\in[n_{\ell}]$,